Hi,
I am doing a question in a assignment where I have to transform the equation to make 'y' the subject with a coefficient of 1. My equation is 29x+4y=11.
I hope this all makes sense, it would be very helpful and I would be very appreciate if you could please solve this problem for me with the working showing so I will know how to do it next time I get a question like this

Answer by ewatrrr(24785)

make 'y' the subject ... solve for y
29x+4y=11
4y = 11-29x
y = (11-29x)/4 Important to use Parenthesis